OpenStudy (anonymous):

Kind of stuck on this one! Please check what I have so far? :) Thanks! Solve x2 + 4x + 8 = 0 x= -2 + 4i

OpenStudy (anonymous):

if \(-2+4i\) is right then the other answer is \(-2-4i\)

OpenStudy (anonymous):

My answer was x= -2 + 4i

OpenStudy (anonymous):

\[x^2+4x+8=0\] \[x^2+4x=-8\] \[(x+2)^2=-8+4=-4\] \[x+2=\pm\sqrt{-4}\] \[x+2=\pm2i\] \[x=-2\pm2i\]

OpenStudy (anonymous):

i think you forgot the square root and write \(\sqrt{-4}=-4i\) but it should be \(\sqrt{-4}=\sqrt{4}\sqrt{-1}=2i\)

OpenStudy (anonymous):

quadratic formula will give the same result, but in this case it is easier to complete the square, because the middle term \(4x\) has an even coefficient

OpenStudy (anonymous):

My mistake, i didn't even catch onto that error! wow thank you for making sense of this!!!!!!!